Grade 9Algebra

if log{(a+b)/2}=1/2(loga + log b)
prove that a=b
please provide me with a solution

Sahil Raj
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
log{(a+b)/2}=1/2 log(ab)
=log(ab)1/2
(a+b)/2= (ab)1/2
a+b=2(ab)1/2
(va)2 +(vb )2-2vavb=0
 
(va -vb)2=0
 
va-vb=0
 
va=vb
 
a=b